OpenStudy (anonymous):

What is the relative maximum and minimum of the function? f(x) = 2x2 + 28x - 8 A. Minimum Value: -106 Range y > -7 B. Minimum Value: -106 Range y > -106 C. Minimum Value: 7 Range y > 7 D. Minimum Value: -7 Range y > -7

OpenStudy (anonymous):

Its not A but i think its D:)

OpenStudy (anonymous):

@skullpatrol

OpenStudy (agent0smith):

find the vertex, use \[\large x = \frac{ -b }{ 2a }\]

OpenStudy (anonymous):

(0.28,0) ?

OpenStudy (anonymous):

Or im not sure

OpenStudy (agent0smith):

One step at a time. For\[\large f(x) = ax^2 + bx +c\]use\[\large x = \frac{ -b }{ 2a }\]

OpenStudy (anonymous):

x= -28/2*2?

OpenStudy (anonymous):

-7?

OpenStudy (agent0smith):

Correct. Now plug that into the equation, to get the minimum.

OpenStudy (anonymous):

f(x)= 2(-7)^2+28(-7)-8 ?

OpenStudy (agent0smith):

yes

OpenStudy (anonymous):

Okay well this is the part im bad at umm -7*-7 is 49 right?

OpenStudy (agent0smith):

Yes, just follow order of operations.

OpenStudy (anonymous):

98....idk how to do this part +28(-7)-8

OpenStudy (anonymous):

-106?

OpenStudy (agent0smith):

Just simplify one thing at a time 2(-7)^2+28(-7)-8 2*49 - 196 - 8 98 - 196 - 8 -106

OpenStudy (anonymous):

Ok so my answer would be A right? I put A and i got it wrong

OpenStudy (agent0smith):

No, -7 has nothing to do with the range. The range is y > the minimum value, which is -106

OpenStudy (anonymous):

Oh so it's B ?

OpenStudy (agent0smith):

Yes

OpenStudy (anonymous):

Thanks so much ! :)

OpenStudy (agent0smith):

You're welcome! But, you said you got it wrong already??!

OpenStudy (anonymous):

Yeah, but im correcting my test to retake it and get a better grade /.\

OpenStudy (agent0smith):

Oh okay